Fermi National Laboratory
 

Accelerator Update for September 13, 2006 - September 15, 2006

 

Wednesday September 13

The day shift began with Tevatron (TeV) experts investigating a sector E48 kicker problem, with Operations monitoring a Recycler stash, with the Detector at Fermilab (CDF) and D-Zero (D0) waiting for the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, and with NuMI and MiniBooNE taking beam.

Operations put the TeV into shot setup at 10:30 AM. After the first antiproton injection into the TeV, the Recycler low level RF system seized.

Operations began preparing the Linac, Booster, Main Injector and other areas for accesses at 12:16 PM.

Operations established store 4952 at 1:52 PM with an average initial luminosity of 16.8E31.

Operations established beam to the Booster at 5:09 PM.

MI access was completed by 7:30 PM.

At 7:41 PM, the first stacking pulse of beam quenched the TeV at F11. The P1 transfer line power supply I:V714 was found in the wrong mode. The subsequent pulses induced a vacuum burst in the P1 line. Vacuum was slow to recover; vacuum technicians were called in.

Operations had established beam to NuMI and MiniBooNE by 7:45 PM.

Operations began preparing D0, CDF, and the TeV A-E enclosures for accesses at 10:12 PM.

CDF was out of access by 11:26 PM.

At 11:44 PM, the TeV F sector accesses were canceled. The vacuum got better.

Thursday September 14

The midnight shift began with Operations monitoring the stash and the stack.

At 12:49 AM, a Linac low energy RF station (LRF3) failed. The investigation and repair lasted the rest of the shift.

D0 came out of access at 4:43 AM.

Operations put the TeV through a dry squeeze at 5:05 AM.

Operations established beam to MiniBooNE and NuMI at 8:05 AM.

Pbar resumed stacking at 9:06 AM.

Operations prepared CDF for access at 9:51 AM.

TeV experts began a study period at 2:27 PM.

The Recycler began stashing at 3:16 PM.

The Meson MTest experiment began taking beam at 4:06 PM.

Operations put the TeV into shot setup at 5:30 PM.

Operations established store 4957 at 7:40 PM with an average initial luminosity of 96.04E30.

Friday September 15

The midnight shift began with Operations monitoring store 4957 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations continued to build up the Recycler stash throughout the owl shift.

The Plans for Friday and the Weekend

The plans for today are to keep store 4957 colliding until the end of day shift and then allow TeV experts to conduct end-of-store studies. D0 will take a two hour access. The weekend will consist of stacking, stashing, and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/15/2006    email Fermilab